# Impacts of axion cooling on the direct detection of supernova axions

### v1

@@ 3.3 Total energy carried away by neutrinos and axions
Figure 6: The total energy carried away by axions and neutrinos. The axion total energy is shown in the solid curves and the neutrino total energy is shown in the broken lines. The red curves represent our results and the black curves represent the results from the aNNaNN model in ref. [16].
@@
Long-term simulations for PNS cooling coupled with the axion emission have been performed in refs. [e.g. 10, 13, 14, 15, 16]. In figure 6, we show the result from a recent simulation in ref. [16]. One can find that, as axions carry away the PNS binding energy, the total neutrino energy becomes suppressed. It is notable that, whereas the simulation setups such as the numerical codes and the progenitor models are different for each model, the difference of $E_{a}$ falls within a factor of two.
@@ 4 Conclusion
In our simulations, the progenitor mass is fixed to $9.6M_{\odot}$ because the star explodes without artificial treatments even in the spherically-symmetric geometry. However, the axion luminosity can be dependent on the progenitor models because it is sensitive to the core temperature, as reported in refs. [26, 91]. In order to obtain successful explosions for heavier stars, one should perform multi-dimensional simulations. Recently, axisymmetric supernova simulations coupled with axions (or axion-like particles) have been performed [17, 92], but these models are limited to the accretion phase at $t_{\mathrm{pb}}\lesssim 1$ s. It is desirable to perform long-term multi-dimensional simulations coupled with the axion cooling for a wide range of progenitors to investigate the detectability of supernova axions from each star.
It should be noted that current supernova models suffer from significant uncertainties in microphysics [e.g. 93, 94, 95, 96]. For example, our simulations do not consider neutrino collective oscillations, which can have qualitative impacts on the supernova explosion mechanism [e.g. 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107]. Also, the effects of different nuclear equations of state [e.g. 108, 109] and neutrino transport schemes [e.g. 110, 46, 111, 47] should be investigated. Better understanding of the standard neutrino-heating explosion mechanism will lead to a more accurate prediction of the axion emission from supernova events. At the same time, as discussed in Section 2, treatments of the axion production rates require further investigations.

### v2

@@ 3.3 Total energy carried away by neutrinos and axions
Figure 6: The total energy carried away by axions and neutrinos. The axion total energy is shown in the solid curves and the neutrino total energy is shown in the broken lines. The red curves represent our results and the black curves represent the results from the aNNaNN model in ref. [2021PhRvD.104j3012F].
@@
Long-term simulations for PNS cooling coupled with the axion emission have been performed in refs. [e.g. PhysRevD.39.1020, 1997PhRvD..56.2419K, 2016PhRvD..94h5012F, 2019JCAP...10..016C, 2021PhRvD.104j3012F]. In figure 6, we show the result from a recent simulation in ref. [2021PhRvD.104j3012F]. One can find that, as axions carry away the PNS binding energy, the total neutrino energy becomes suppressed. It is notable that, whereas the simulation setups such as the numerical codes and the progenitor models are different for each model, the difference of $E_{a}$ falls within a factor of two.
@@ 4 Conclusion
In our simulations, the progenitor mass is fixed to $9.6M_{\odot}$ because the star explodes without artificial treatments even in the spherically-symmetric geometry. However, the axion luminosity can be dependent on the progenitor models because it is sensitive to the core temperature, as reported in refs. [2022PhRvD.105f3028C, 2025arXiv250309005T]. In order to obtain successful explosions for heavier stars, one should perform multi-dimensional simulations. Recently, axisymmetric supernova simulations coupled with axions (or axion-like particles) have been performed [PhysRevD.106.063019, 2023PhRvD.108f3027M], but these models are limited to the accretion phase at $t_{\mathrm{pb}}\lesssim 1$ s. It is desirable to perform long-term multi-dimensional simulations coupled with the axion cooling for a wide range of progenitors to investigate the detectability of supernova axions from each star.
It should be noted that current supernova models suffer from significant uncertainties in microphysics [e.g. Kotake06, Janka12, Burrows21, Yamada2024]. For example, our simulations do not consider neutrino collective oscillations, which can have qualitative impacts on the supernova explosion mechanism [e.g. 2019PhRvD.100d3004A, 2019ApJ...886..139N, 2021PhRvD.103f3033A, 2021PhRvD.104h3025N, 2023PhRvD.108l3024L, 2023PhRvD.107h3016X, Nagakura2023, Ehring2023a, Ehring2023b, 2024PhRvD.109b3012A, 2025PASJ...77...L9M, 2025arXiv250722985F]. Also, the effects of different nuclear equations of state [e.g. Sumiyoshi2020, 2021PrPNP.12003879B] and neutrino transport schemes [e.g. 2009ApJ...698.1174L, 2011PThPh.125.1255S, 2012ApJS..199...17S, 2013PhRvD..87j3004C] should be investigated. Better understanding of the standard neutrino-heating explosion mechanism will lead to a more accurate prediction of the axion emission from supernova events. At the same time, as discussed in Section 2, treatments of the axion production rates require further investigations.